a collection of 39 nickles and dimes is worth $3.10. Find the number of nickels and dimes.
Answer by CubeyThePenguin(3113)

n = number of nickels
d = number of dimes

n + d = 39
5n + 10d = 310

(eliminate d)

10n + 10d = 390
-5n - 10d = -310
----------------------
5n = 80 ---> n = 16

---> 16 nickels, 23 dimes